OpenStudy (anonymous):

ln(4x-2) = ln(8x) I keep getting -1/2 but the answer is no solution, I dont get why?

jagr2713 (jagr2713):

whats in

zepdrix (zepdrix):

\[\Large\rm 4x-2=8x\]Solving for x gives us,\[\Large\rm x=-\frac{1}{2}\]But you need to plug it back into the equation to make sure the solution works.

zepdrix (zepdrix):

\[\Large\rm \ln(-4)\]Uh oh! :O

hartnn (hartnn):

log of negative numbers is not defined in real domain.

OpenStudy (anonymous):

oh ok thanks, i get it now

OpenStudy (isaiah.feynman):

Good!

OpenStudy (anonymous):

completely forgot about negative logs lol

OpenStudy (isaiah.feynman):

No such thing as negative logs my friend.

hartnn (hartnn):

there is, but in complex domain

OpenStudy (isaiah.feynman):

I must study more advanced mathematics then